Suppose nominal gdp starts at 1000 and grows each quarter by 0.65% . What is the level of gap after 7 years?
1
Expert's answer
2019-05-17T12:09:18-0400

according to the compound interest formula:

gdp after 7 years = 1000*(1+0.65/100) 28 =1198

where 28 = 4 quarters * 7 years

1198 - 1000 = 198

the level of gap after 7 years = 198